# Wiener-Hammerstein LTI Linear-Time Invariant

### v1

@@ 2 Problem Formulation
Let $\mathcal{X}=\{\mathbf{x}^{(m)}\}_{j=1}^{M}$ denote a dataset of source audio signals, where each $\mathbf{x}\in\mathbb{R}^{L}$ is assumed to be drawn from a prior distribution $p_{x}$. Additionally, let $\mathcal{Z}_{0}=\{\mathbf{z}_{0}^{(n)}\}_{n=1}^{N}$ represent an independent dataset of signals also drawn from the same distribution $p_{x}$. Importantly, the dataset $\mathcal{Z}_{0}$ is unobserved; instead, we are provided with a dataset $\mathcal{Y}=\{\mathbf{y}^{(n)}\}_{n=1}^{N}$ of effected measurements. Each observed signal $\mathbf{y}\in\mathbb{R}^{L}$ is assumed to be generated from a corresponding clean source signal $\mathbf{z}_{0}$ through an unknown distortion process, described by a function
@@ 3.1 Proposed Diffusion-Based Approach
[…]
@@
The key challenge in optimizing Eq. (4) is evaluating the expectation, which requires drawing samples from the posterior distribution $p(\mathcal{Z}_{0}|\mathcal{Y})$. Since the dataset $\mathcal{Y}$ consists of $N$ distorted signals $\mathbf{y}^{(n)}$, the algorithm requires estimating the corresponding source input signals $\mathbf{z}_{0}^{(n)}$ for each $n\in\{1,\dots,N\}$. For clarity, we omit the index $n$ in the following discussion, but the following procedure must be repeated for each observation in the dataset.
@@ 3.1.1 Diffusion models for posterior sampling
where time $\tau$ evolves backward from $T$ to 00. Since the score function $\nabla_{\mathbf{z}_{\tau}}\log p(\mathbf{z}_{\tau})$ is generally intractable, it is approximated using a time-conditional deep neural network $s_{\theta}(\mathbf{z}_{\tau},\tau)\approx\nabla_{\mathbf{z}_{\tau}}\log p(\mathbf{z}_{\tau})$, trained via denoising score matching [33] on the available dataset of clean signals $\mathcal{X}$. We adopt the same diffusion parameterization as in previous works [25, 24] and refer to [25] for a more detailed introduction.
@@ 3.1.2 E-Step
$\{\tau_{1},\dots,\tau_{k-1},\tau_{k},\dots,\tau_{K}\},\quad\text{with}\quad\tau_{1}=T\quad\text{and}\quad\tau_{K}\approx 0.$
@@
Following [32], we further approximate $p(\mathbf{z}_{0}|\mathbf{z}_{\tau})$ as a Dirac delta located at the one-step denoised estimate: $p(\mathbf{z}_{0}|\mathbf{z}_{\tau})\approx\delta(\hat{\mathbf{z}}_{0}(\mathbf{z}_{\tau}))$. This results in the following one-sample Monte Carlo estimate of the expectation from Eq. (4):
@@ 3.1.3 M-Step
[…]
@@ 3.2 Adversarial Approach
The second approach, sketched in Fig. 1(b), is inspired by Gated Convolutional Networks [34] and was proposed for guitar amplifier modeling by Wright et al. [13]. It consists of the following optimization objective [29]:
[…]
@@
$\quad+\mathbb{E}_{\mathbf{x}\sim p_{x}}\big{[}-\max(0,1+D(\hat{f}(\mathbf{x};\psi);\phi))\big{]}\Bigg{\}}$ (14)
@@ 4.1 Black-Box Operator Models
Black-box models are data-driven and do not require explicit knowledge of the underlying system in their design. Deep neural networks, as universal function approximators, are commonly used to approximate the input-output behavior of an effect. In this study, we focus on two specific neural architectures that have proven succesful at modeling audio effects: a Gated Convolution Network (GCN) and S4, a state-space model [36].
@@ 4.2 Wiener-Hammerstein Model
One of the most widely used approaches for modeling nonlinear systems are Wiener-Hammerstein models, which represent the system as a serial connection of a Linear-Time Invariant (LTI) block, followed by a static nonlinearity, and another LTI block [42, 43, 44, 45, 40]. Unlike black-box models, gray-box approaches like W-H structures offer better interpretability of the learned transfer function, allowing a detailed analysis of each block after the optimization process. In the context of unsupervised optimization, we hypothesize that a more targeted model with a constrained parameter space, such as a W-H model, can be particularly beneficial. The model’s inductive biases, derived from prior knowledge of the unknown operator, can potentially help guide the optimization, improving stability and robustness. However, this advantage may come at the cost of reduced expressivity, as the model structure may limit its ability to fit highly complex systems that cannot be predicted with the utilized parametric structure.
@@ 5.1 Data
Following [13], we utilize the fourth subset of the IDMT-SMT Guitar dataset [48], which comprises 64 short musical excerpts spanning various styles and tempos. While the dataset includes recordings from two different guitars, we exclusively use those from the Career SG guitar. Each recording is processed with three distinct distortion effects of increasing severity: ‘Clean Distortion,’ ‘Light Distortion,’ and ‘Heavy Distortion.’ All the material is sampled at 44.1 kHz.
Our experiments require the following distinct, non-overlapping dataset splits: two unpaired datasets representing the input and output distributions, denoted as $\mathcal{X}$ and $\mathcal{Y}$, respectively, and a paired test set ($\mathcal{X}_{\text{test}}$, $\mathcal{Y}_{\text{test}}$) for evaluation. Of the 37 min of available recordings, we allocate 16 min for $\mathcal{X}$, and 16 min for $\mathcal{Y}$. Additionally, we construct three reduced versions of $\mathcal{Y}$ containing 18 s, 1 min, and 4 min of distorted recordings. The remaining 5 min are used for the test set, which is segmented into chunks of 666\,6s, resulting in 50 segments. Our dataset split differs from the one used in [13], as this configuration enables a more systematic evaluation of both approaches. Consequently, our results may not be directly comparable to those reported in [13].
@@ 5.2 Experimental Details: Diffusion-Based Approach
Our diffusion-based method closely follows the configuration of previous work [25], using the same hyperparameter choices unless otherwise specified. We employ a waveform-domain diffusion model with the CQT-based architecture proposed in [22], training it on 6-second audio segments. Preliminary studies indicated that this architecture generalizes better than STFT-based or waveform-based alternatives, performing better for this task. The model was pre-trained for 80k iterations using the EGDB dataset [49], followed by 14k iterations of training on the dataset split $\mathcal{X}$, with a batch size of 4. The training phase (excluding pre-training) took 70 min [^3] All computation times here were measured on an NVIDIA H200 GPU.. While the impact of pre-training remains to be formally evaluated, preliminary experiments suggest that it may have only a minimal effect.
We use a consistent hyperparameter configuration for all the operator optimization experiments, setting $T=101$ steps, which corresponds to both the reverse diffusion discretization and the number of EM iterations. The likelihood scaling parameter $\zeta$ in Eq. (5) is defined following [25] as a function of $\tau$ controlled by $\tilde{\zeta}=0.2$. As the cost function $C(\cdot,\cdot)$, we employ a $\ell_{2}$-norm in a compressed STFT representation [23, 24], using a compression factor of 0.5. Each M-step includes 20 operator optimization steps. We employ the AdamW optimizer with random batches of 4 examples per iteration, a learning rate of 1e-3, momentum parameters ($\beta_{1}=0.9$, $\beta_{2}=0.99$), and a weight decay of 1e-2.
@@ 5.4 Experimental Details: Supervised Baseline
In addition to the unsupervised methods that are the focus of this study, we incorporate a supervised baseline to serve as a performance upper bound for the unsupervised approaches. This baseline is trained using the dataset $\mathcal{Y}$, with reduced size when applicable, along with the paired input signal, which is unavailable in the unsupervised setting. The supervised operator models were trained for 5k iterations using the Adam optimizer with a learning rate of 1e-4 and 6-second-long audio segments.
@@ 5.5 Evaluation
The operator $g$ is the AFXx-Rep trained encoder, $\cdot$ denotes the dot product, the norm used is the $\ell_{2}$ norm, and $\epsilon$ is a small constant for numerical stability.
@@ 5.6 Results
The experiments with the diffusion-based method outperform the adversarial ones in the ‘Clean’ and ‘Light’ distortion scenarios when only 18s or 1 min of effected recordings are available, and achieves a comparable performance when 4 or 16 min are used. In terms of AFx-Rep dist. (Fig.3a,b) and $\ell_{1}\!\log\mathrm{MSS}$ (Fig.3d,e), the adversarial approach yields slightly better results when the full dataset is available. Conversely, the diffusion-based method achieves consistently lower errors on the linear $\ell_{1}\!\,\mathrm{MSS}$ metric (Fig.3g,h) These differences are potentially attributed to the alignment between each method training objective and the evaluation metrics. The adversarial model was trained using log-scaled mel-spectrogram features, which are more closely related to the $\ell_{1}\!\log\mathrm{MSS}$ metric. In contrast, the diffusion-based model was optimized using magnitude-compressed STFT features, which may be more aligned with linear-scale $\ell_{1}\,\mathrm{MSS}$, though the correspondence is not exact.
@@
We do not observe a clear trend regarding the most suitable black-box model architecture: both GCN (solid lines) and S4 (dashed lines) yield comparable performance when used with either of the unsupervised methods. While one may outperform the other in specific cases, the difference is not consistent across settings. Within the diffusion-based framework, a comparison between black-box and grey-box operators shows that the W-H model performs on par with both the GCN and S4 models (see Fig. 3), and slightly better in some cases. Attempts to train the W-H model using the adversarial approach were unsuccessful, and the corresponding results are therefore not included in the figure. These findings suggest that diffusion-based methods offer stable performance across different operator architectures, whereas the adversarial approach appears more sensitive to the choice of model.
@@ 6 Conclusions
This study addressed unsupervised operator estimation, with experiments on guitar distortion modeling, comparing diffusion-based and adversarial approaches. While adversarial methods performed well under heavy distortion, they lack consistency across scenarios. In contrast, diffusion-based methods show strong robustness to data scarcity and operator choice, making them a more reliable option overall. Additionally, diffusion methods offer the benefit of jointly estimating the clean guitar signal, although reconstruction quality has not been explored in this work. Finally, while all operator models are capable of real-time processing [1], their computational demands remain an open question for future investigation.

### v2

@@ 2 Problem Formulation
Let $\mathcal{X}=\{\mathbf{x}^{(m)}\}_{m=1}^{M}$ denote a dataset of source audio signals, where each $\mathbf{x}\in\mathbb{R}^{L}$ is assumed to be drawn from a prior distribution $p_{x}$. Additionally, let $\mathcal{Z}_{0}=\{\mathbf{z}_{0}^{(n)}\}_{n=1}^{N}$ represent an independent dataset of signals also drawn from the same distribution $p_{x}$. Importantly, the dataset $\mathcal{Z}_{0}$ is unobserved; instead, we are provided with a dataset $\mathcal{Y}=\{\mathbf{y}^{(n)}\}_{n=1}^{N}$ of effected measurements. Each observed signal $\mathbf{y}\in\mathbb{R}^{L}$ is assumed to be generated from a corresponding clean source signal $\mathbf{z}_{0}$ through an unknown distortion process, described by a function
@@ 3.1 Proposed Diffusion-Based Approach
$\underset{\psi}{\mathrm{max}}\,\mathbb{E}_{\mathcal{Z}_{0}\sim p(\mathcal{Z}_{0}|\mathcal{Y})}\log p(\mathcal{Y}|\mathcal{Z}_{0};\psi)\,$, (4)
@@
The key challenge in optimizing Eq. (4) is evaluating the expectation, which requires drawing samples from the posterior distribution $p(\mathcal{Z}_{0}|\mathcal{Y})$. Since the dataset $\mathcal{Y}$ consists of $N$ distorted signals $\mathbf{y}^{(n)}$, the algorithm requires estimating the corresponding source input signals $\mathbf{z}_{0}^{(n)}$ for each $n\in\{1,\dots,N\}$. For clarity, we omit the index $i$ in the following discussion, but the following procedure must be repeated for each observation in the dataset.
@@ 3.1.1 Diffusion models for posterior sampling
where time $\tau$ evolves backward from $T$ to 0. Since the score function $\nabla_{\mathbf{z}_{\tau}}\!\log p(\mathbf{z}_{\tau})$ is generally intractable, it is approximated using a time-conditional deep neural network $s_{\theta}(\mathbf{z}_{\tau},\tau)\approx\nabla_{\mathbf{z}_{\tau}}\!\log p(\mathbf{z}_{\tau})$, trained via denoising score matching [33] on the available dataset of clean signals $\mathcal{X}$. We adopt the same diffusion parameterization as in previous works [25, 24] and refer to [25] for a more detailed introduction.
@@ 3.1.2 E-Step
$\{\tau_{1},\dots,\tau_{k-1},\tau_{k},\dots,\tau_{K}\},\qquad\text{with}\qquad\tau_{1}=T\qquad\text{and}\qquad\tau_{K}\approx 0.$
@@
Following [32], we further approximate $p(\mathbf{z}_{0}|\mathbf{z}_{\tau})$ as a Dirac delta located at the one-step denoised estimate: $p(\mathbf{z}_{0}|\mathbf{z}_{\tau})\approx\delta(\hat{\mathbf{z}}_{0}(\mathbf{z}_{\tau},\tau))$. This results in the following one-sample Monte Carlo estimate of the expectation from Eq. (4):
@@ 3.1.3 M-Step
$\psi\leftarrow\underset{\psi}{\mathrm{arg\,min}}\sum_{n=1}^{N}\mathcal{C}\left(\mathbf{y}^{(n)}\!,\,\hat{f}(\hat{\mathbf{z}}_{0}(\mathbf{z}_{\tau_{k}}^{(n)},\tau);\psi)\right),$ (12)
@@ 3.2 Adversarial Approach
The second approach, sketched in Fig. 1(b), is inspired by Generative Adversarial Networks [34] and was proposed for guitar amplifier modeling by Wright et al. [13]. It consists of the following optimization objective [29]:
$\psi=\underset{\psi}{\text{arg min}}\,\mathcal{D}(p_{y},\hat{p}_{y}^{\psi}),$ (13)
@@
$\qquad+\mathbb{E}_{\mathbf{x}\sim p_{x}}\big[-\max(0,1+D(\hat{f}(\mathbf{x};\psi);\phi))\big]\Bigg\}$ (14)
@@ 4.1 Black-Box Operator Models
Black-box models are data-driven and do not require explicit knowledge of the underlying system in their design. Deep neural networks, as universal function approximators, are commonly used to approximate the input-output behavior of an effect. In this study, we focus on two specific neural architectures that have proven successful at modeling audio effects: a Gated Convolution Network (GCN) and S4, a state-space model [36].
@@ 4.2 Wiener-Hammerstein Model
One of the most widely used approaches for modeling nonlinear systems is the Wiener-Hammerstein model, which represents the system as a serial connection of a Linear-Time Invariant (LTI) block, followed by a static nonlinearity, and another LTI block [42, 43, 44, 40]. Unlike black-box models, gray-box approaches such as W-H structures offer better interpretability of the learned transfer function, allowing a detailed analysis of each block after the optimization process. In the context of unsupervised optimization, we hypothesize that a more targeted model with a constrained parameter space, such as a W-H model, can be particularly beneficial. The model’s inductive biases, derived from prior knowledge of the unknown operator, can potentially help guide the optimization, improving stability and robustness. However, this advantage may come at the cost of reduced expressivity, as the model structure may limit its ability to fit highly complex systems that cannot be predicted with the utilized parametric structure.
@@ 5.1 Data
Following [13], we utilize the fourth subset of the IDMT-SMT Guitar dataset [47], which comprises 64 short musical excerpts spanning various styles and tempos. While the dataset includes recordings from two different guitars, we exclusively use those from the Career SG guitar. Each recording is processed with three distinct distortion effects of increasing severity: ‘Clean Distortion,’ ‘Light Distortion,’ and ‘Heavy Distortion.’ The distortion effects were applied using commercial audio plugins, consistent with the procedure in [13]. All the material is sampled at 44.1 kHz.
Our experiments require the following distinct, non-overlapping dataset splits: two unpaired datasets representing the input and output distributions, denoted as $\mathcal{X}$ and $\mathcal{Y}$, respectively, and a paired test set ($\mathcal{X}_{\text{test}}$, $\mathcal{Y}_{\text{test}}$) for evaluation. Of the 37 min of available recordings, we allocate 16 min for $\mathcal{X}$, and 16 min for $\mathcal{Y}$. Additionally, we construct three reduced versions of $\mathcal{Y}$ containing 18 s, 1 min, and 4 min of distorted recordings. The remaining 5 min are used for the test set, which is segmented into chunks of 6 s, resulting in 50 segments. Our dataset split differs from the one used in [13], as this configuration enables a more systematic evaluation of both approaches. Consequently, our results may not be directly comparable to those reported in [13].
@@ 5.2 Experimental Details: Diffusion-Based Approach
Our diffusion-based method closely follows the configuration of previous work [25], using the same hyperparameter choices unless otherwise specified. The diffusion model operates in the waveform domain, while its score network is built on an architecture based on the Constant-Q Transform (CQT), as proposed in [22]. This design leverages the invertibility and differentiability of the CQT to introduce inductive biases tailored to musical signals, while retaining the flexibility of waveform-domain modeling. The model is trained on 6-second audio segments. The model was pre-trained for 80k iterations using the EGDB dataset [48], followed by 14k iterations of training on the dataset split $\mathcal{X}$, with a batch size of 4. The training phase (excluding pre-training) took 70 min [^3] All computation times here were measured on an NVIDIA H200 GPU.. While the impact of pre-training remains to be formally evaluated, preliminary experiments suggest that it may have only a minimal effect.
We use a consistent hyperparameter configuration for all the operator optimization experiments, setting $T=101$ steps, which corresponds to both the reverse diffusion discretization and the number of EM iterations. The likelihood scaling parameter $\zeta$ in Eq. (5) is defined following [25] as a function of $\tau$ controlled by $\tilde{\zeta}=0.2$. As the cost function $C(\cdot,\cdot)$, we employ a $\ell_{2}$-norm in a compressed STFT representation [23, 24], using a compression factor of 0.5. This compression equalizes the spectral energy distribution and enhances high-frequency content, which typically has lower energy but is perceptually important. Each M-step includes 20 operator optimization steps. We employ the AdamW optimizer with random batches of 4 examples per iteration, a learning rate of 0.001, momentum parameters ($\beta_{1}=0.9$, $\beta_{2}=0.99$), and a weight decay of 0.01.
@@ 5.3 Experimental Details: Adversarial Approach
We acknowledge the extensive body of work aimed at improving the stability of adversarial training [49, 50], and recognize that some of these techniques could be relevant to our setting, particularly in imbalanced scenarios. However, exploring and properly implementing these approaches is considered beyond the scope of the present work. Therefore, we limit ourselves to the configuration proposed in [13].
@@ 5.4 Experimental Details: Supervised Baseline
In addition to the unsupervised methods that are the focus of this study, we incorporate a supervised baseline to serve as an upper performance bound for the unsupervised approaches. This baseline is trained using the dataset $\mathcal{Y}$, with reduced size when applicable, along with the paired input signal, which is unavailable in the unsupervised setting. The supervised operator models were trained for 5k iterations using the Adam optimizer with a learning rate of 0.0001 and 6-s-long audio segments.
@@ 5.5 Evaluation
The operator $g$ is the AFx-Rep trained encoder, $\cdot$ denotes the dot product, the norm used is the $\ell_{2}$ norm, and $\epsilon$ is a small constant for numerical stability.
@@ 5.6 Results
The experiments with the diffusion-based method outperform the adversarial ones in the ‘Clean’ and ‘Light’ distortion scenarios when only 18 s or 1 min of effected recordings are available, and achieves a comparable performance when 4 or 16 min are used. In terms of AFx-Rep dist. (Fig. 3a,b) and $\ell_{1}\!\log\mathrm{MSS}$ (Fig. 3d,e), the adversarial approach yields slightly better results when the full dataset is available. Conversely, the diffusion-based method achieves consistently lower errors on the linear $\ell_{1}\!\,\mathrm{MSS}$ metric (Fig. 3g,h). These differences are potentially attributed to the alignment between each method’s training objective and the evaluation metrics. The adversarial model was trained using log-scaled mel-spectrogram features, which are more closely related to the $\ell_{1}\!\log\mathrm{MSS}$ metric. In contrast, the diffusion-based model was optimized using magnitude-compressed STFT features, which may be more aligned with linear-scale $\ell_{1}\,\mathrm{MSS}$, though the correspondence is not exact.
@@
We do not observe a clear trend regarding the most suitable black-box model architecture: both GCN and S4 yield comparable performance when used with either of the unsupervised methods. While one may outperform the other in specific cases, the difference is not consistent across settings. Within the diffusion-based framework, a comparison between black-box and grey-box operators shows that the W-H model performs on par with both the GCN and S4 models (see Fig. 3), and slightly better in some cases. Attempts to train the W-H model using the adversarial approach were unsuccessful, and the corresponding results are therefore not included in the figure. These findings suggest that diffusion-based methods offer stable performance across different operator architectures, whereas the adversarial approach appears more sensitive to the choice of model.
@@ 6 Conclusions
This study addressed unsupervised operator estimation, with experiments on guitar distortion modeling, comparing diffusion-based and adversarial approaches. While adversarial methods performed well under heavy distortion, they lack consistency across scenarios. In contrast, diffusion-based methods show strong robustness to data scarcity and operator choice, making them a more reliable option overall. Additionally, diffusion methods offer the benefit of jointly estimating the clean guitar signal, although reconstruction quality has not been explored in this work.
Both approaches—diffusion-based and adversarial—require substantial training time and computational resources, which can limit their accessibility for applications that require training with low-power or time constraints. A key limitation of the diffusion-based method is the need for a separately pre-trained model on clean guitar signals, and obtaining such dry data can be challenging depending on the context. Future work should investigate how performance varies with different amounts of available dry data and assess the trade-offs between robustness and computational efficiency. Although all evaluated models are technically capable of real-time operation [1], the practical computational demands remain an open area for further study. Finally, although this study focused specifically on distortion effects, we believe the proposed framework could extend to other nonlinear audio effects, such as dynamic range compression or modulation effects, given the generality of the black-box operators used in training.